# L1: Controlling How Long A Reasoning Model Thinks With Reinforcement Learning

### v1

@@ Appendix A Results
Method Mean Error (%$\downarrow$ RMSE Error (%$\downarrow$
@@
Given the surprisingly strong performance of LCPO, particularly at lower token budgets, a natural question arises regarding whether the model has memorized solutions, suggesting potential train-test leakage. To address this concern, we specifically evaluate on the AIME2025 dataset, an exam released after the training of all evaluated models. Results in Table 3 demonstrate that the performance drop for LCPO is substantially smaller ($\approx 2\%$) compared to other models such as Agentica-24K (-12.9%), DeepSeek-R1-1.5B (-7.5%), and even Llama-3.3-70B (-23.5%). These results provide strong evidence that despite using shorter CoT lengths, our model demonstrates generalization rather than relying solely on memorization.
A.3 Parallel vs Sequential Scaling
@@
Figure 10 compares how sequential scaling (increasing CoT length) and parallel scaling (Majority Voting) vary in performance for the same incurred cost (total tokens generated). There are a few interesting observations from the graph: First, parallel scaling is almost always worse than sequential scaling, when compared at the same token budget. Second, parallel scaling can provide significant improvements, to all reasoning models (eg, more than 10% in some cases). Third, while Qwen-2.5-1.5B-Math, barely benefits from majority voting, L1-Max (1024 tokens), despite using similar number of generation tokens in each task, benefits significantly. This is noteworthy, since for many cases, it is desirable to have low token lengths, because of both reduced latency, and avoiding the quadratic cost of generation. In such cases, LCPO-Max clearly severs as a better alternative to parallel scaling of non-reasoning models. Finally, LCPO-Max (3600 tokens) matches the performance of Agentica-4K baseline across parallel scaling. We do note a small dip in the last point, but this is because only 1 seed was used for parallel scaling, and as such the difference is not statistically significant.
A.4 Qualitative Examples
[…]
@@ A.4 Qualitative Examples
[…]
@@
In Figures 11 and 12, we show two examples of L1-Exact solving the same problem but with different requested generation lengths: 512 and 3600 tokens. While the model with 512 tokens is unable to solve the problem correctly, the model with longer generation arrives at the correct answer. We refer interested readers to our website, which has more such interesting examples.

### v2

@@ Appendix A Results
[…]
@@
Figure 10: RMSE in Length of generated chains across different datasets. LCPO exhibits low RMSE for math reasoning datasets, maintaining high precision in length control. OOD datasets exhibit higher RMSE, particularly MMLU, where longer chains are less necessary.
@@
Figure 12: Hard Budget Violation Rates across different datasets. L1 maintains relatively higher, but acceptable violation rates of 9% on average across different token budgets.
Figure 13: Soft Budget Violation Rates across different datasets. L1 maintains less than 3% budget violation rate on across different token budgets.
Further, we also report both soft- and hard-violation rates in Figure 13 and Figure 13. Notably, while soft-violation rates are very low ($\leq 3\%$), hard-violation rates are relatively higher (9% on average across different token budgets). The violation rates are particularly higher at lower token budgets, as the model prioritizes performance over length control, and generating correct solutions with shorter CoTs is relatively more difficult.
@@
Given the surprisingly strong performance of LCPO, particularly at lower token budgets, a natural question arises regarding whether the model has memorized solutions, suggesting potential train-test leakage. To address this concern, we specifically evaluate on the AIME2025 dataset, an exam released after the training of all evaluated models. Results in Table 3 demonstrate that the performance drop for LCPO is substantially smaller ($\approx 2\%$) compared to other models such as DeepScaleR-24K (-12.9%), DeepSeek-R1-1.5B (-7.5%), and even Llama-3.3-70B (-23.5%). These results provide strong evidence that despite using shorter CoT lengths, our model demonstrates generalization rather than relying solely on memorization.
A.3 Aha Moment in L1’s training
Figure 14 shows the training logs of L1-Exact. During the first 300 RL steps, we observe that the model prioritizes improving its average solve rate while the minimum response length continues to decrease. This suggests that the model initially focuses on correctly generating solutions rather than strictly adhering to length constraints. However, at approximately 300 RL steps, we observe a significant phase transition in the training dynamics. Specifically, the reward and token adherence score ($\text{reward}-\text{solve\_mean}$) begin to increase rapidly, while the minimum response length exhibits a sharp drop, eventually converging at around 100 tokens (the lower bound of $n_{gold}$). This distinct transition point represents when the model effectively learns the concept of token adherence and subsequently continues to refine this capability throughout further training.
Figure 14: Training Logs for LCPO-Exact. We notice an Aha Moment, where the model starts adhering to token constraint after 300 RL steps, as visible in peaking solve rate, and sharp drop in minimum token length.
A.4 Does L1 generalize to varying sampling parameters
Temperature Mean Deviation Solve Rate (%) Avg. Token Length
0.0 6.0% 46.6% 1778
0.3 6.3% 46.6% 1782
0.6 6.2% 46.5% 1778
1.0 5.6% 46.4% 1765
Table 4: Effect of sampling temperature on length control and accuracy for L1-Exact. Despite changing the temperature, the model shows similar errors in length control and similar Pass@1 solve rates throughout all temperature settings.
We evaluate L1-Exact’s robustness to different sampling parameters by varying the temperature during inference. Table 4 presents results across temperature values ranging from 0.0 to 1.0. We observe that the model maintains consistent performance in terms of both solve rate and length control accuracy. Specifically, the mean deviation from target length remains stable (5.6-6.3%) across all temperature settings, with marginal differences in token length and solve rate. These results indicate that L1’s length control capabilities are robust to sampling parameters, allowing flexible deployment across different inference settings without compromising performance or adherence to length constraints. Further, given that we average across four target lengths $n_{gold}\in\{512,1024,2048,3600\}$, the observed average token lengths (1765-1782) closely match the ideal average of 1796 tokens.
A.5 Prior Reasoning Models do not follow length constraints in Prompts
Requested Length 512 1024 2048 3600
Generated Length 5491 6072 5846 6421
Table 5: Average length generated by DeepScaleR-24K for different requested lengths. The model does not follow the user-specified length constraint.
A natural question is whether existing reasoning models can follow length constraints through simple prompting without specialized training. To investigate this, we explicitly prompted DeepScaleR-24K to generate responses of specific token lengths (512, 1024, 2048, and 3600) using clear prompts like ”Think for exactly N tokens.” As shown in Table 5, the model consistently generates outputs close to 6000 tokens regardless of the requested length, demonstrating a complete insensitivity to length instructions.
A.6 Supervised Fine-Tuning does not learn length control
We further compare how our proposed RL recipe compares to SFT. For training models with supervised finetune, we first generate responses without any length control, then measure their token lengths and re-prompt with the measured token counts to create instruction-response pairs with explicit length requirements. For training the models, we perform a hyperparameter search over Learning Rates: 1e-6, 1e-5, 1e-4 and Batch Sizes: 256, 512. We use validation loss for early stopping.
After training with this SFT data, we evaluated adherence by prompting for specific lengths. The results are presented in Table 6. The results highlights that the model still failed to follow length constraints, producing very long outputs regardless of the requested target.
Requested Tokens 512 1024 2048 4096
Actual Tokens 21388 22749 21426 20903
Table 6: SFT-only model does not learn to follow token-length constraints despite being trained on relabeled data with explicit length requirements.
These results, along with Table 5, indicate that offline methods such as SFT are ineffective at inducing precise length control in reasoning models. We hypothesize two reasons: (1) the narrow distribution of output lengths for a given question leads the model to ignore length instructions, and (2) without an online length-sensitive reward, the model does not acquire the capability to adjust its reasoning to meet target budgets.
A.7 Ablating Reward Functions in LCPO
We evaluate how different objective functions affect the performance and token adherence of LCPO. Specifically, we ablate Equation 2 using the following variants:
  1. 1.
    L1-Max Single Objective: During the training of L1-Max, we originally employed both objectives from Eq. 2 and Eq. 1. In this variant, we exclusively use the former.
  2. 2.
    L1-Max Addition: This variant implements a simple addition of correctness and length penalty objectives instead of multiplication. The objective function is: $r(y,y_{gold},n_{gold})=\mathbb{I}(y=y_{gold})-\alpha\cdot\bigl|n_{gold}-n_{y}\bigr|$.
  3. 3.
    L1-Max Sigmoid: This variant applies a sigmoid function to penalize length violations using the objective: $r(y,y_{gold},n_{gold})=\mathbb{I}(y=y_{gold})\cdot\sigma(\alpha\cdot(n_{gold}-n_{y}))$, where $\sigma(x)=\frac{1}{1+e^{-x}}$ is the sigmoid function.
We train each variant for 120 RL steps, initializing from the same L1-Exact checkpoint. Figure 17 presents the budget violation rates across these variants. While all variants exhibit similar violation rates across different token budgets, L1-Max Addition achieves a 0% budget violation rate consistently. This perfect adherence, however, occurs because the model collapses to generating extremely short CoTs, which trivially satisfies the maximum length constraint which is undesirable behavior.
Figure 15 demonstrates this effect by showing performance versus token length for all variants. LCPO-Max Addition performs substantially worse than the other three variants. For clarity, we exclude this variant in Figure 16. The remaining three variants perform similarly without any statistically significant differences. We ultimately select the standard L1-Max for all our experiments due to its simplicity and the potential advantages of its dual objective approach.
Figure 15: Performance of different Reward Functions (RFs) used to train L1-Max.
Figure 16: Performance of different Reward Functions (RFs) used to train L1-Max.
Figure 17: Budget Violation Rates for different Reward Functions used to train LCPO-Max.
A.8 Parallel vs Sequential Scaling
@@
Figure 18 compares how sequential scaling (increasing CoT length) and parallel scaling (Majority Voting) vary in performance for the same incurred cost (total tokens generated). There are a few interesting observations from the graph: First, parallel scaling is almost always worse than sequential scaling, when compared at the same token budget. Second, parallel scaling can provide significant improvements, to all reasoning models (eg, more than 10% in some cases). Third, while Qwen-2.5-1.5B-Math, barely benefits from majority voting, L1-Max (1024 tokens), despite using similar number of generation tokens in each task, benefits significantly. This is noteworthy, since for many cases, it is desirable to have low token lengths, because of both reduced latency, and avoiding the quadratic cost of generation. In such cases, LCPO-Max clearly severs as a better alternative to parallel scaling of non-reasoning models. Finally, LCPO-Max (3600 tokens) matches the performance of DeepScaleR-4K baseline across parallel scaling. We do note a small dip in the last point, but this is because only 1 seed was used for parallel scaling, and as such the difference is not statistically significant.
A.9 LCPO scales to larger reasoning models
Figure 19: Scaling LCPO to a 7B reasoning model. We apply LCPO to DeepSeek-R1-Distill-7B reasoning model and show strong length control and performance on math reasoning tasks.
We further scaled the LCPO recipe to DeepSeek-R1-Distill-7B model. Results are presented in Figure 19. We observe the same trends as with the 1.5B model: high controllability, low budget-violation rates, and strong performance relative to length-controlled baselines across budgets. These results highlight the effectiveness and potential of LCPO to scale to even larger reasoning models.
A.10 Qualitative Examples
Example Question Find the largest possible real part of $(75+117i)z+\frac{96+144i}{z}]$ where $z$ is a complex number with $|z|=4$. Let’s think step by step and output the final answer within boxed{}. Think for 512 tokens. Correct Answer $\boxed{540}$ Correctness p Incorrect Tokens Requested 512 Model Response Thinking Process Okay, so I have this problem: I need to find the largest possible real part of the expression $(75+117i)z+(96+144i)/z$, where $z$ is a complex number with $|z|=4$. Hmm, let’s break this down. First, since $|z|=4$, $z$ can be written as $4e^{i\theta}$ for some angle $\theta$. That might help simplify things. Let me denote $z=4e^{i\theta}$. Then, $1/z=(1/4)e^{-i\theta}$ because $|z|=4$. So, the expression becomes $(75+117i)\cdot 4e^{i\theta}+(96+144i)\cdot(1/4)e^{-i\theta}$. Simplifying, that’s $4(75+117i)e^{i\theta}+(24+36i)e^{-i\theta}$. Now, I need to find the real part of this expression. Let’s denote $\text{Re}[(75+117i)z+(96+144i)/z]$. Breaking it down, $\text{Re}[(75+117i)z]+\text{Re}[(96+144i)/z]$. Since $z=4e^{i\theta}$, $\text{Re}[(75+117i)\cdot 4e^{i\theta}]+\text{Re}[(96+144i)\cdot(1/4)e^{-i\theta}$. This becomes $4\text{Re}[(75+117i)e^{i\theta}]+(1/4)\text{Re}[(96+144i)e^{-i\theta}]$. Now, let’s compute $\text{Re}[(75+117i)e^{i\theta}]$. This is $75\cos\theta+117\sin\theta$. Similarly, $\text{Re}[(96+144i)e^{-i\theta}]$ is $96\cos\theta-144\sin\theta$. So, putting it all together, the real part is $4(75\cos\theta+117\sin\theta)+(1/4)(96\cos\theta-144\sin\theta)$. Simplifying, that’s $300\cos\theta+468\sin\theta+24\cos\theta-36\sin\theta$. Combining like terms, we get $(300+24)\cos\theta+(468-36)\sin\theta=324\cos\theta+432\sin\theta$. Now, we need to find the maximum of $324\cos\theta+432\sin\theta$. This can be written as $R\cos(\theta-\phi)$, where $R=\sqrt{324^{2}+432^{2}}$. Calculating $R$: $324^{2}=104,976$, $432^{2}=186,624$. Sum is $391,600$. $R=\sqrt{391,600}\approx 625.6$. Thus, the maximum real part is approximately 625.6. The largest possible real part is $\boxed{625.6}$.
@@ A.10 Qualitative Examples
Example Question Find the largest possible real part of $(75+117i)z+\frac{96+144i}{z}]$ where $z$ is a complex number with $|z|=4$. Let’s think step by step and output the final answer within $boxed\{\}$. Think for 3600 tokens. Correct Answer $\boxed{540}$ Correctness Correct Tokens Requested 3600 Model Response Thinking Process (not shown to user) Okay, so I have this problem: I need to find the largest possible real part of the expression (75 + 117i)z + (96 + 144i)/z, where z is a complex number with $|z|=4$. Hmm, that seems a bit involved, but let’s break it down step by step. First, let me understand what is given. We have a complex number z with magnitude 4, so z can be written in the form z = $4e^{i\theta}$, where $\theta$ is the argument of z. That might be useful because it allows us to express z in terms of its angle, which could simplify the expression. The expression we need to consider is $(75+117i)z+\frac{96+144i}{z}$. We are asked to find the largest possible real part of this expression. So, let’s denote the expression as $E=(75+117i)z+\frac{96+144i}{z}$. [Reasoning process truncated for brevity …] Then, adding $104,976+186,624=291,600$. $\sqrt{291,600}=\sqrt{2916*100}=\sqrt{2916}*10$. Now, $\sqrt{2916}$: let’s see, $54^{2}=2916$, because $50^{2}=2500$, $54^{2}=2916$. So, $\sqrt{2916}=54$, thus $\sqrt{2916*100}=54*10=540$. So, that seems correct. Therefore, the maximum value of $\text{Re}(E)$ is 540. Therefore, the largest possible real part is 540. Given the expression $(75+117i)z+\frac{96+144i}{z}$ where $z$ is a complex number with $|z|=4$, we need to find the largest possible real part of this expression. First, we express $z$ as $4e^{i\theta}$ since $|z|=4$. Then, we substitute $z$ and $\frac{1}{z}$ into the expression: $(75+117i)z+\frac{96+144i}{z}$ [Output truncated for brevity …] Thus, the largest possible real part is: $\boxed{540}$
@@
In Figures 20 and 21, we show two examples of L1-Exact solving the same problem but with different requested generation lengths: 512 and 3600 tokens. While the model with 512 tokens is unable to solve the problem correctly, the model with longer generation arrives at the correct answer. We refer interested readers to our website, which has more such interesting examples.